# Policy Learning with a Language Bottleneck

### v1

@@ 5 SelectSay
As a proof-of-concept, we first consider SelectSay, a simple collaborative game introduced in Hu and Sadigh (2023). Here, a speaker can see a hidden set of five balls including two blue ones and must help a listener find the two blue balls by communicating with them only via numbers, see Figure 3. In principle, the speaker and listener could converge on any bijective convention between messages and balls, and this is what we find with RL agents (multiple optimal policies). Humans, on the other hand, would prefer the intuitive mapping 1 $\rightarrow$ 1, 2 $\rightarrow$ 2, etc.  that we call the human-interpretable policy. Hu and Sadigh (2023) showed that regularizing the listener’s policy with the instruction I should select the same number as my partner, successfully guided the two RL agents to the human-interpretable policy. We investigate whether PLLB could generate a similar rule and achieve a similar interpretability on its own.
@@
Figure 4: SelectSay. a) Bottleneck agents learn as fast as TabularQ and InstructRL and faster than agents using adversarial rules Adversarial. b) Unlike TabularQ, they learn human-interpretable policies; this without relying on an external instruction like InstructRL. c) When faced with speakers enforcing a non-human-interpretable policy, Bottleneck converges faster than InstructRL.
@@ 5.2 PLLB can learn counter-intuitive policies
A potential confound in the preceding experiment is the possibility that PLLB helped converge to the human-interpretable policy because this is the only behavior an LM is able to describe. To test this hypothesis we fixed the speaker to use a counter-intuitive policy using a random mapping between balls and messages (e.g.,  1 $\rightarrow$ 3, 5 $\rightarrow$ 2, etc.). Here, a listener following the human-interpretable policy would fail. In Figure 4c we observe that while InstructRL and Adversarial, both regularized by misaligned rules (the human-interpretable one for InstructRL, another random permutation for Adversarial), are able to eventually adapt to the fixed speaker, they converge at a slower rate than both TabularQ (no rule) and Bottleneck. Overall, Bottleneck is the only algorithm able to flexibly adapt to a fixed speaker policy while converging on the human-like policy when the speaker is also learning.
@@ 6 Maze
Task Overview. We randomly generate different 7x7 mazes with the gym-maze code base. [^5] https://github.com/MattChanTK/gym-maze Agents are rewarded when they achieve the goal, by the inverse of the number of steps they needed to reach it as an incentive to learn the optimal solving path. We introduce additional structure to all mazes by coloring cells for which the optimal action is south (red) or north then east (blue) with 50% probability (or they are left blank). Therefore, agents that leverage color information may better generalize to new mazes where the coloring structure is preserved, but the optimal action sequence differs.
The first iteration of PLLB correspond to training a standard RL algorithm to obtain $\pi_{1}$, in our case a tabular Q-learning agent (TabularQ), since there are no rules yet. After the agent observes at least two solved mazes, we run gen_rule by prompting a llama-2-70b-chat LM with contrastive episodes and a task description. This gives us a linguistic rule $\mathcal{L}_{1}$ (e.g., Upon observing RED, take SOUTH) that we can use to update the policy. We first induce the regularizing policy $\pi_{\mathcal{L}_{1}}$ with rule $\mathcal{L}_{1}$ by obtaining a probability distribution over the 4 actions from the LM, and then run the RL algorithm for 5 episodes to obtain the new policy $\pi_{2}$ (procedure described in Section 3). We repeat these steps every 5 episodes of interactions with the environment.
Baselines include the base tabular Q-learning algorithm without language bottleneck (TabularQ) and a variant of Bottleneck generating rules from reward-randomized episode samples (Adversarial). We additionally evaluate LinearQ, an agent learning a linear model Q-function with an additional feature for cell color. We train LinearQ with a batch update of size 10 and learning rate 0.001, after performing a hyperparameter sweep, and note that we found it generally took significantly longer to converge than the other methods.
@@ 6.1 PLLB learns more generalizable policies
Figure 5a shows that learning a valid rule (Bottleneck) does not increase learning speed over TabularQ (already quite fast), but using a corrupted rule (Adversarial) or learning from linear features (LinearQ) slows down learning. We next replace the maze by another 7x7 maze sharing the same underlying color semantics, but with a different optimal action sequences. For fair comparison, we start with the fully converged policy for each method (requiring 100 episodes for LinearQ). Figure 5b shows that Bottleneck leverages the learned rule and adapts to the new maze more effectively than both TabularQ and a fully converged LinearQ agent. While TabularQ cannot generalize because it does not perceive colors, LinearQ does generalize faster at first but converges slower.
We find that the generated rules indeed improve over time to better capture the underlying structure of the maze (e.g., if I observe BLUE, then take the NORTH action; see other examples in Appendix  B.1). Across all trials, $100\%$ of the final rules mention the red $\rightarrow$ south rule and $60\%$ uncovered the more complex blue $\rightarrow$ north-then-east rule.
@@ 6.2 PLLB learns adaptable policies
In Figure 5c, we reproduce this same experiment on a maze with a different underlying structure (red now indicates west while blue indicates east then south). Although the rule Bottleneck learned does not apply anymore, it can still adapt faster than the baselines. The language bottleneck can indeed quickly capture the new structure of the task and regularize learning to steer adaptation. Here LinearQ seems to have overfit to the first maze and struggled to adapt. We find that all trials end with rules capturing the new mapping (100% red $\rightarrow$ west, 50% blue $\rightarrow$ east-then-south). Overall, these results show that the language bottleneck supports human-like cognitive functions by finding a tradeoff between the efficiency of TabularQ and the generalization capabilities LinearQ while remaining more adaptable.
@@ 6.3 PLLB is more interpretable and inter-operable
Figure 6: Participants provided rules from PLLB solve new mazes faster than those given either visual or no aid (control), left. They also self-reported linguistic feedback as more useful than visual (right).
@@
Participants using PLLB rules solve the new maze with fewer steps than others and find this aid more useful than the visual one in average (statistically significant at level 0.05 with two-sided Mann-Whitney test after Bonferroni correction, see Figure 6). Participants indeed found it harder to extract the relevant information from the visual aid, which contained a great deal of non-transferable information (optimal actions in all non-colored cells). PLLB rules focus on the useful and transferable information learned by the previous agent, which is much more readily usable by humans.
Overall, the results in Maze demonstrate the ability of PLLB to train agents that are more generalizable and adaptable (cognitive use) as well as more interpretable by, and inter-operable with humans (communicative use).
@@ 7.1 Task Overview
For both tasks, we consider two settings: From Scratch, where the listener is initialized using the original pre-trained model weights at every episode, and Continual Training, where the listener is continuously updated over time using descriptions by the speaker. We therefore split our datasets into separate held-out sets for training, early stopping, selecting samples for gen_rule, and evaluation.
@@ 7.3 PLLB can collaborate with human listeners
Do human listeners also benefit from image descriptions provided by speakers following Bottleneck rules? We next conduct a human subject study for the Builder task, where we replace the listener with study participants at the end of one iteration of PLLB, and evaluate how accurately and quickly they can reconstruct target images from a held-out test set. We recruit 20 crowdworkers on Prolific to reconstruct 5 images using descriptions from a speaker following a rule $\mathcal{L}$ generated with (Bottleneck), and 5 images using the original (Baseline) descriptions. We provided participants an interface that included a drawing canvas and buttons controlling shape types and colors (Section A.1), allowing us to use our reward function of Levenshtein similarity between user actions and the underlying action sequences for a target image.
@@ 7.4 PLLB collaborates with human speakers
The descriptions of workers not provided a rule are more diverse and less focused e.g., A barbed wire is an uncomfortable stop for a bird. With PLLB rules on the other hand (e.g., Describe the bird’s coloration accurately), they generate more specific descriptions: Red and black bird on barbed wire. Bright red chest, red at top of head, black wings and beak for the same target image (see other example rules in Appendix Table 4). PLLB agents can easily transmit what they learned from experience (the rule) to humans who can then benefit from that knowledge (cultural transmission, communicative use).
@@ 8 Limitations & Discussion
By inferring the rule explaining its best past behaviors, PLLB becomes more interpretable by and inter-operable with humans. When it uses a rule to condition its policy updates, it steers its own behavior to become more interpretable and generalizable. In our current implementation, rule generation remains limited: we require converting agent histories into representations the LM can handle (e.g., text or images), which may prevent the use of long-horizon sensorimotor trajectories (e.g., robotic systems). Rule generation can also fail when the LM overfits or abstracts too much (e.g., in Birds we found the rule: “Use descriptive language that conveys the mood, such as the serenity of a snowy day or the freedom of flight”). Advances in multi-modal models and better approaches for selecting contrastive examples for gen_rule might help mitigate these issues.
An interesting direction for future work is to apply PLLB to tasks with complex reward functions that include hard-to-articulate human preferences. In applications like image captioning for accessibility (Nie et al., 2020) and personalization of language models (Li et al., 2023a), linguistic rules might improve transparency and predictability even in cases where users find it difficult to describe their own goals.
Finally, we could extend PLLB to generate other kinds of linguistic information: not only abstract rules, but also explanations, goals, or learning strategies. Such information could then be shared with others (cultural transmission) or used by update functions in new ways: mapping generated language to state or action abstractions, goals, reward functions, or even learning strategies. This opens the door for more human-like AI agents leveraging the deep interactions between language and decision-making (Colas et al., 2022).
@@ Acknowledgements
We would like to acknowledge support from NSF Awards #2006388, #2125511, #2238240 and #2212310, AFOSR YIP, ONR #N000142112298, and DARPA. Megha Srivastava was additionally supported by an IBM PhD Fellowship. Cédric Colas received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065949.
@@ Appendix A Additional Methods
A.1 Maze

### v2

@@ 5 SaySelect
As a proof-of-concept, we first consider SaySelect, a simple collaborative game introduced in Hu and Sadigh (2023). Here, a speaker can see a hidden set of five balls including two blue ones and must help a listener find the two blue balls by communicating with them only via numbers, see Figure 3. In principle, the speaker and listener could converge on any bijective convention between messages and balls, and this is what we find with RL agents (multiple optimal policies). Humans, on the other hand, would prefer the intuitive mapping 1 $\rightarrow$ 1, 2 $\rightarrow$ 2, etc.  that we call the human-interpretable policy. Hu and Sadigh (2023) showed that regularizing the listener’s policy with the instruction I should select the same number as my partner, successfully guided the two RL agents to the human-interpretable policy. We investigate whether PLLB could generate a similar rule and achieve a similar interpretability on its own.
@@
Figure 4: SaySelect. a) Bottleneck agents learn as fast as TabularQ and InstructRL and faster than agents using adversarial rules Adversarial. b) Unlike TabularQ, they learn human-interpretable policies; this without relying on an external instruction like InstructRL. c) When faced with speakers enforcing a non-human-interpretable policy, Bottleneck converges faster than InstructRL.
@@ 5.2 PLLB can learn counter-intuitive policies
A potential confound in the preceding experiment is the possibility that PLLB helps converge to a human-interpretable policy because this is the only behavior an LM can describe. To test this hypothesis, we fixed the speaker to use a counter-intuitive policy using a random mapping between balls and messages (e.g.,  1 $\rightarrow$ 3, 5 $\rightarrow$ 2, etc.). Here, the human-interpretable policy is mis-aligned, and would lead the listener to fail or learn slower. In Figure 4c we observe that while InstructRL and Adversarial, both regularized by misaligned rules (the human-interpretable one for InstructRL, another random permutation for Adversarial), are able to eventually adapt to the fixed speaker, they converge at a slower rate than both TabularQ (no rule) and Bottleneck. This effect was more pronounced at higher values of $\lambda$. Overall, Bottleneck is the only method that flexibly adapts to a fixed speaker policy while also converging on the human-like policy when the speaker is learning, without human supervision.
@@ 6 Maze
Task Overview. We randomly generate different 7x7 mazes with the gym-maze code base. [^5] https://github.com/MattChanTK/gym-maze Agents receive as reward the inverse of the number of steps they needed to reach the goal as an incentive to learn the optimal solution. Although all mazes are random, we introduction additional structures by coloring cells for which the optimal action is south (red) or north then east (blue) with probability 50% (or they are left blank). Therefore, agents that leverage color information can better generalize to new mazes where the coloring is preserved, even if the optimal action sequence differs.
The first iteration of PLLB corresponds to training a standard RL algorithm to obtain $\pi_{1}$, in our case a tabular Q-learning agent (TabularQ). After the agent observes at two solved mazes, we run gen_rule by prompting a llama-2-70b-chat LM with contrastive episodes and a task description. This gives us a linguistic rule $\mathcal{L}_{1}$ (e.g., Upon observing RED, take SOUTH) that we can use to update the policy. We first induce the regularizing policy $\pi_{\mathcal{L}_{1}}$ with rule $\mathcal{L}_{1}$ by obtaining a probability distribution over the 4 actions from the LM, and then run the RL algorithm for 5 episodes to obtain the new policy $\pi_{2}$ (procedure described in Section 3). We repeat these steps every 5 episodes of interactions with the environment.
Baselines include the tabular Q-learning algorithm without language bottleneck (TabularQ) and a variant of Bottleneck generating rules from reward-randomized episode samples (Adversarial). We also evaluate LinearQ, an agent learning a linear model Q-function with an additional feature for cell color. We train LinearQ with a batch size of 10 and learning rate 0.001, after performing a hyperparameter sweep. We found LinearQ took significantly longer to converge than the other methods, likely due to the increased state representational complexity.
@@ 6.1 PLLB learns more generalizable policies
Figure 5a shows that, in the Standard setting, learning a valid rule (Bottleneck) does not increase learning speed of a single maze over TabularQ. However, using a corrupted rule (Adversarial) or learning from linear features (LinearQ) slows down learning. Furthermore, Bottleneck outperforms TabularQ with respect to generalization: when we switch to a new 7x7 maze sharing the same underlying color semantics, but with a different optimal action sequences. Bottleneck leverages the learned rule and generalized to the new maze more effectively than all other agents (Figure 5b). For a a fair comparison, we start with the fully converged policy for each method (requiring 100 episodes for LinearQ). While TabularQ cannot generalize because it does not perceive colors, LinearQ does generalize faster at first, but converges slower.
The generated rules improve over time to better capture the underlying structure of the maze (e.g., if I observe BLUE, then take the NORTH action; see other examples in Appendix  B.1). Across all trials, $100\%$ of the final rules mention the red $\rightarrow$ south rule and $60\%$ uncovered the more complex blue $\rightarrow$ north-then-east rule. Finally, we find that a policy that does not update Q-values (that is, purely implemented via the LM) achieves a success rate $0\%$ within the same time limit, even when using the final rules generated from PLLB. This highlights the importance of leveraging experience that cannot be expressed with language, which text-based reasoning agents cannot do.
@@ 6.2 PLLB learns adaptable policies
In Figure 5c, we reproduce this same experiment on a maze with a different underlying structure (red now indicates west while blue indicates east then south). Although the rule Bottleneck learned does not apply anymore, it can still adapt faster than the baselines. We find that all trials end with rules capturing the new mapping (100% red $\rightarrow$ west, 50% blue $\rightarrow$ east-then-south). Meanwhile, LinearQ overfits to the first maze and struggle to adapt. Overall, our results show that PLLB supports human-like cognitive functions by finding a tradeoff between the efficiency of TabularQ and the generalization capabilities LinearQ, while remaining more adaptable.
@@ 6.3 PLLB is more interpretable and inter-operable
Figure 6: Participants provided rules from PLLB solve new mazes faster than those given either visual or no aid (control), left. They also self-reported linguistic feedback as more useful than visual (right). * marks statistical significance at level 0.05 with two-sided Mann-Whitney test after Bonferroni correction,
@@
Participants using PLLB rules solve the new maze with fewer steps than others and find this aid significantly more useful than the visual one in average (see Figure 6). Participants indeed found it harder to extract the relevant information from the visual aid, which contained a great deal of non-transferable information (optimal actions in all non-colored cells). PLLB rules focus on the useful and transferable information learned by the previous agent, which is much more readily usable by humans. Overall, the results in Maze demonstrate the ability of PLLB to train agents that are more generalizable and adaptable (cognitive use) as well as more interpretable by, and inter-operable with humans (communicative use).
@@ 7.1 Task Overview
For both tasks, we consider two settings: From Scratch, where the listener is initialized using the original pre-trained model weights at every episode, and Continual Training, where the listener is continuously updated over time using descriptions by the speaker. We create separate held-out splits for training, early stopping, selecting samples for gen_rule, and evaluation.
@@ 7.3 PLLB can collaborate with human listeners
Do human listeners benefit from image descriptions provided by speakers following Bottleneck rules? We conduct a human subject study for the Builder task, where we replace the listener with study participants at the end of one iteration of PLLB, and evaluate how quickly they can reconstruct target images from a held-out test set. We recruit 20 crowdworkers on Prolific to reconstruct 5 images using descriptions from a speaker following a rule $\mathcal{L}$ generated with (Bottleneck), and 5 images using the original (Baseline) descriptions. We provided participants an interface that included a drawing canvas and buttons controlling shape types and colors (Section A.2), allowing us to use our reward function of Levenshtein similarity between (sorted) user actions and the underlying action sequences for a target image.
@@ 7.4 PLLB collaborates with human speakers
The descriptions from users not provided a rule are more diverse and less focused e.g., A barbed wire is an uncomfortable stop for a bird. With PLLB rules on the other hand (e.g., Describe the bird’s coloration accurately), they generate more specific descriptions: Red and black bird on barbed wire. Bright red chest, red at top of head, black wings and beak for the same target image (see other example rules in Appendix Table 4). PLLB agents can easily transmit what they learned from experience (the rule) to humans who can then benefit from that knowledge (cultural transmission, communicative use).
8 Robot Grasp Planning
Figure 11: Results in Grasp. a) Bottleneck agents learn faster than the non-linguistic RL Baseline agents, as well as Adversarial and an Oracle zero-shot prompt of a VLM without any internal policy update. b) When faced with unseen bag designs and locations, Bottleneck agents generalize and learn faster than both Baseline (which over-relies on image details) and Oracle, ( c) shows the fixed linguistic rule used). d) Examples grasps, shaded by reward, and generated rules $\mathcal{L}$ for pairs of contrasting episodes.
Can PLLB rules improve policy learning in embodied tasks with complex action spaces such as robot planning? Existing works on language model-based agents (e.g., ReAct (Yao et al., 2022) or Reflexion (Shinn et al., 2024)) and robot policies (e.g., Code as Policies) leverage pre-trained models for action generation, but cannot be applied in domains where a (non-LM-based) policy must be updated using real-world reinforcement learning signals.
Consider training a robot to lift a paper bag. The most intuitive grasp would be at the bag’s handles; indeed, a zero-shot Oracle prompt of the molmo-7b multi-modal language model outputs grasp keypoints on bag handles (Figure 11). However, if the bag is too heavy, a robot should adapt to less common grasps (e.g. by the side) to avoid tearing. Because VLMs are not grounded with actual physical quantities, such adaptation is impossible without leveraging an external reward (Gao et al., 2024). On the other hand, existing RL-based approaches for robotic manipulation are often inefficient (Kalashnikov et al., 2018). We show that PLLB rules can be used to efficiently learn non-intuitive robot grasping policies.
8.1 Task Overview
In the Grasp task, the AI agent is a Franka Emika Panda Arm robot that needs to lift and move a paper bag filled with unknown objects off of a shelf, as shown in Appendix Figure 15. Due to the complex action space, we follow existing work by decomposing the robot’s policy into (i) perceiving the scene and identifying an appropriate grasp location, and (ii) planning a path to this location (ten Pas et al., 2017; Kalashnikov et al., 2018).
We begin with a fixed keypoint-conditioned path-planning policy which is pre-trained on a dataset of 50 trajectories with 3D-point cloud observations using Sphinx (Sundaresan et al., 2025). This model outputs a trajectory of 7-DoF end-effector poses (x/y/z/yaw/pitch/row/gripper) for the robot to execute given the keypoint.
We use PLLB to train (via RL) a grasp prediction policy that selects a grasp keypoint (a 2-D position) given an image of the environment. We define a successful grasp as one where the robot can lift the bag without it breaking, and so set our reward to be 0 for a successful grasp, -100 if the bag is not grasped at all, and use the normalized downward torque (e.g. -20Nm) applied to the robot’s shoulder joint as a dense reward for unsuccessful grasps. Initially, the agent performs random grasps. After it observes one successful grasp, we run gen_rule with the open-source molmo-7b vision–language model with contrastive images of successful and unsuccessful grasp keypoints. This gives us a linguistic rule $\mathcal{L}$ (e.g. grasp the left side of the bag) that we can use to update the policy. We regularize the initial grasping policy toward the policy $\pi_{\mathcal{L}}$ induced from rule $\mathcal{L}$ by prompting molmo-7b with $\mathcal{L}$, then instructing it to generate random keypoints consistent with the description.
We compare our Bottleneck method with an Adversarial version (swapping high and low reward grasps), a Vanilla RL baseline that only uses image observation data, and an Oracle that always guides the policy with the rule generated from zero-shot prompting molmo-7b (no constrasting examples based on a reward).
8.2 PLLB can learn counter-intuitive policies
As shown in Figure 11, when averaged across 20 random seeds, our Bottleneck converges to higher reward compared to all other methods. Because the Oracle method suggests a rule encouraging the policy to predict key points near the bag’s handles, it requires more episodes to learn that this behavior leads to lower reward. Qualitatively, we observe the generated rules converge towards guiding the robot to grasp at less common locations (e.g., Grasp the bottom of the object, see examples in Appendix Section B.1). This shows that PLLB can help improve over standard RL methods even for learning counter-intuitive policies.
8.3 PLLB learns more generalizable policies
To evaluate generalization, we create an environment set-up with a new bag, location, and set of visual distractions (see Appendix 15). As shown in Figure 11, the rules learned by Bottleneck outperform alternative methods for generalizing to this new environment. Because the Vanilla RL policy can only rely on image features, it is not able to leverage the more generalizable strategy PLLB rules capture. Overall, the results in Grasp show that PLLB can improve policy convergence even in the complex dynamics of embodied tasks like robotic grasping.
@@ 9 Limitations & Discussion
PLLB helps agents becomes more interpretable by and inter-operable with humans. In our current implementation, rule_gen requires converting agent episodes into representations an LM can handle (e.g., text or images), which may prevent the use of long-horizon sensorimotor trajectories. Nevertheless, our Grasp task shows how PLLB can leverage progress in pretrained models for robotic tasks, and future advances in long context modeling can address this limitation. Rule generation can also fail when the LM abstracts too much (e.g., in Birds we found the rule: “Use descriptive language that conveys the mood, such as the serenity of a snowy day or the freedom of flight”). Advances in multi-modal models and better approaches for selecting contrastive examples can also mitigate these issues (Appendix B.2).
An interesting direction for future work is applying PLLB to tasks with complex reward functions that reflect hard-to-articulate human preferences, such as image captioning for accessibility (Nie et al., 2020) and personalized language models (Li et al., 2023a). Linguistic rules could enhance transparency and predictability even when users struggle to express their goals.
Additionally, PLLB could generate not just abstract rules but also explanations, goals, or learning strategies, enabling cultural transmission or novel uses in update functions—such as mapping language to state/action abstractions, goals, or reward functions. This paves the way for more human-like AI agents that creatively integrate language with decision-making (Colas et al., 2022).
@@ Acknowledgements
We would like to acknowledge support from NSF Awards #2006388, #2125511, #2238240 and #2212310, AFOSR YIP, ONR #N000142112298, and DARPA. Research was sponsored by the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. Megha Srivastava was additionally supported by an IBM PhD Fellowship. Cédric Colas received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065949.
@@ Appendix A Additional Methods
A.1 Hyperparameters
  1. 1.
    SaySelect: We use the same default parameters used in InstructRL (Hu and Sadigh, 2023), including setting the regularization strength $\lambda=0.25$. Additionally, each time we invoke gen_rule, we create an ensemble of 3 rules, which we aggregate over when construct the probability distribution over actions during update.
  2. 2.
    Maze: We use the InstructRL objective with a tabular Q-learning agent, but introduce a $\epsilon_{LM}$ parameter that controls whether the regularization strength $\lambda$ is 0 or 1 at each time-step in an episode. Although we did not observe a strong effect on modifying $\epsilon_{LM}$, we did not explore values larger than $\epsilon_{LM}=0.4$ as that led to increased inference cost and experiment latency. Finally, each time we invoke gen_rule we create an ensemble of 4 rules, which we aggregate over when construct the probability distribution over actions during update.
  3. 3.
    Builder: For the listener agent we finetune BART for a maximum of 100 epochs at each iteration of PLLB, employing early stopping on a held-out validation set and using the default arguments provided by the HuggingFace Transformers library (Lewis et al., 2019). Each time we invoke gen_rule, we same 3 rules and select the rule with the highest aggregate likelihood across all tokens. Likewise, for each image description (the speaker’s action), we sample 3 possible descriptions under the given rule and select the one with the highest probability.
  4. 4.
    Birds: For the fine-tuned version of the listener agent, we finetune StableDiffusion for 1000 steps on a separate finetuning dataset. When generating images, we simply sample one image per description, using 10 inference steps and a guidance scale of 7.5.
  5. 5.
    Grasp: For grasp keypoint selection,we train an upper confidence bound agent on a neural contextual bandit using image features as context, with learning rate 0.01, with an $(100,100)$ action space containing integer coordinates of the environment image. During update, we sample 20 coordinates following a given rule from gen_rule, which we use as the probability distribution over actions, setting $\lambda=0.2$ as the regularization strength. To execute the grasp, we use a grasp model built by training Sphinx (Sundaresan et al., 2025) over 50 waypoint trajectories with learning rate 0.0001 for 100 epochs.
A.2 Maze
See Figures 12 and 13.